OpenStudy (anonymous):

what is a three step equation in which the answer is 13?

OpenStudy (anonymous):

lol - another one of these!!

OpenStudy (anonymous):

oh - jetly is gone - id better do it

OpenStudy (anonymous):

ok 6x + 17 = x + 82

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Didn't you post about 1000 of these the other day? ;) \[\Large 7a - 45 = 3a + 7\]

OpenStudy (anonymous):

OK what is a three equation in which the answer is 15?